Determine the percent by mass of each element in CuBr2.
Answer #1

Use 1 mole of the compound

molar mass CuBr2 = 63.55 + (2 x 79.9) = 233.35 g/mol

mass Cu in 1 mole = 63.55 g

% mass Cu = mass Cu / total mass CuBr2 x 100/1
= 63.55 g / 233.35 g x 100
= 27.2 %
Answer #2

Use 1 mole of the compound

molar mass CuBr2 = 63.55 + (2 x 79.9) = 233.35 g/mol

mass Cu in 1 mole = 63.55 g

% mass Cu = mass Cu / total mass CuBr2 x 100/1
= 63.55 g / 233.35 g x 100
= 27.2 %

for Br2=100-27.2=72.8%